OpenStudy (anonymous):

\int\limits \frac{ x^2-1 }{ x^3\sqrt{2x^4-2x^2+1} }

OpenStudy (anonymous):

\[\int\limits \frac{ x^2-1 }{ x^3\sqrt{2x^4-2x^2+1} }dx\]

hartnn (hartnn):

just trying..... whats the derivative of \(\Large \dfrac{2x^4-2x^2+1}{x^4}\) did you get it something like 1/x-1/x^3 ...

hartnn (hartnn):

if yes, then you can put that as u and see what happens

OpenStudy (anonymous):

i started off with a substitution like x^2=t..but then failed

OpenStudy (anonymous):

derivative of the above is ..\[\frac{ 4(x^2-1)}{ x^5 }\]

hartnn (hartnn):

\(\Large \dfrac{2x^4-2x^2+1}{x^4} =2 - \dfrac{2}{x^2} + x^{-4} \\ d= 4/x -4/x^3 = 4 (x^2-1)/x^3 \)

OpenStudy (anonymous):

umm ..yeahh ..i think i messed that up with u/v rule ..!

hartnn (hartnn):

so we have the form \(\Large (1/4)\int \dfrac{d(f(x))}{\sqrt {f(x)}}\) form

hartnn (hartnn):

no we don't ...i guess my method/trial failed....

OpenStudy (anonymous):

yeahh ..i have the answer ...do u need that ? would dat be any help ?

hartnn (hartnn):

i have the answer too.

OpenStudy (anonymous):

umm do u have this .?\[\frac{ \sqrt{2x^4-2x^2+1} }{ 2x^2 } + c\]

hartnn (hartnn):

yup

OpenStudy (anonymous):

how did u get that ?

hartnn (hartnn):

@myininaya @RadEn @ganeshie8 @mathmale might wanna have a look...

OpenStudy (anonymous):

@ganeshie8 ??

ganeshie8 (ganeshie8):

\(\large \int\frac{ x^2-1 }{ x^3\sqrt{2x^4-2x^2+1} }dx\) divide top and bottom wid x^2 \(\large \int\frac{ 1-\frac{1}{x^2} }{ x^3\sqrt{2-\frac{2}{x^2}+\frac{1}{x^4}} }dx\) sub u = 1/x^2 du = -2/x^3 dx \(\large \int\frac{ 1-u }{\sqrt{2-2u+u^2} } (-\frac{du}{2})\)

ganeshie8 (ganeshie8):

somplete the square in the bottom, and sub \(u-1 = \tan\theta\)

ganeshie8 (ganeshie8):

*complete

hartnn (hartnn):

aah! one more substitution of u = 1/x^2 did the trick! good catch @ganeshie8

ganeshie8 (ganeshie8):

:)

ganeshie8 (ganeshie8):

humm reverse trig sub can be tricky, if we want the given answer in terms of x's

OpenStudy (anonymous):

amazing ....:!!!! thank you all .!

ganeshie8 (ganeshie8):

np :)